QUESTION 21
A p value is the probability of obtaining the value of the test statistic or more extreme, if the null hypothesis is true.

a.   True

b.   False


QUESTION 18
Type I error occurs when

a.   you reject a false null hypothesis

b.   you reject a true null hypothesis

c.   you fail to reject a false null hypothesis

d.   you fail to reject a true null hypothesis


QUESTION 19
Type II error occurs when

a.   you reject a false null hypothesis

b.   you reject a true null hypothesis

c.   you fail to reject a false null hypothesis

d.   you fail to reject a true null hypothesis

QUESTION 20
Power refers to the probability that

a.   you reject a false null hypothesis

b.   you reject a true null hypothesis

c.   you fail to reject a false null hypothesis

d.   you fail to reject a true null hypothesis
Answer #1

TOPIC:P-value,Type-I error,Type-II error,power of a test.- Question - 21. The answer is as True. Explanation : - The is defined as p-value of a hypothesis testing, a p-value = p. of

= question - 18h The answer is → y You fail to reject the a false null hypothesis, * Explanation - In hypothesis testing theNow, B = P Not rejecting I to when He is actually false] So, Power = (1-P), So, Power = 1-P ( Not rejecting Ho, when to is 1-